OpenStudy (anonymous):

What is the next number in the pattern: 3, 6, 7, 14, 15, 30, 31, 62, 63, …

Parth (parthkohli):

This pattern looks like this - we have a number, then we have its double, then we have the next, then the double

Parth (parthkohli):

next number would be the double of 63

OpenStudy (anonymous):

126

OpenStudy (anonymous):

3+3=6,7 7+7=14,15 15+15=30,31 31+31=62,63 63+63=126

Parth (parthkohli):

@drpick Why did she get a medal?

OpenStudy (anonymous):

cause IMA BOSS!!! lol

Parth (parthkohli):

I asked something

OpenStudy (anonymous):

What is the 8th number in the pattern: 20, 12, 4, -4, …?

OpenStudy (anonymous):

yeah XD

Parth (parthkohli):

This is an arithmetic sequence

OpenStudy (anonymous):

20-8=12 12-8=4 4-8=-4 -4-8=-12 -12-8=-20 -20-8=-28 -28-8=-36<------------ 8th number in the pattern

Parth (parthkohli):

Delete it ^

OpenStudy (anonymous):

???

Parth (parthkohli):

Delete it... because people aren't allowed to be given answers

OpenStudy (anonymous):

No people aren't allowed to give JUST answers with out a direct explanation, I clearly showed him how Ive gotten the answer I have...

Parth (parthkohli):

You're late. You shouldn't have given the answer.

Parth (parthkohli):

@aussy123 people like drpick love the answers. they don't wanna see the explanation.

OpenStudy (anonymous):

Yeah i just dont want the answer...i want to know how to do it SMH

Parth (parthkohli):

I told you how to do it, yet you want the answers.

OpenStudy (anonymous):

well your kinda boring to explain

Parth (parthkohli):

Users like you make the site look bad.

OpenStudy (anonymous):

and you still get mad because i don't give you a medal

Parth (parthkohli):

No. I don't want a medal. I just want you to take a medal away from her.

OpenStudy (anonymous):

why???

Parth (parthkohli):

Because people who ask for answers and who give answers are not welcome here.

Parth (parthkohli):

http://openstudy.com/code-of-conduct Read this. Your excuses wouldn't be allowed here.

Parth (parthkohli):

@Tomas.A would you explain? That's what amby's are made for :)

OpenStudy (anonymous):

ParthKohli says truth.

OpenStudy (anonymous):

I will encourage and guide those needing help, and not just give them an answer.....I didnt just give him the answers I clearly worked it out for him..

OpenStudy (anonymous):

yeah what an epic fail SMH

Parth (parthkohli):

@aussy123 If you really want to help, you will guide the asker and won't give the answer

OpenStudy (anonymous):

he needs to work them out himself especially when someone else is trying to help him, and you just pop there and ruin all his hard job because it's much harder to guide person than solve yourself.

Parth (parthkohli):

The biggest challenge of the site is to explain the answer fiends and the answer snipers of how the site really works

OpenStudy (anonymous):

Dude ParthKohli you happy now you have the medal...so now you can stop harassing me over helping someone...and I gave steps and everything I just didnt write it out in words... Many people on here work it out the same way

Parth (parthkohli):

I have the medal from Tomas, not the asker. and I don't care about medals, get that? I want the site to be a better place. @TuringTest would like to speak more on the topic

OpenStudy (anonymous):

@aussy123 im sorry for getting you into this problem

OpenStudy (anonymous):

its nothing, because it clearly shows that I showed the work ive done and I clearly didnt just give the answers...as long as you understand, thats all that matters

OpenStudy (anonymous):

thanks thats why i gave you the medal

Parth (parthkohli):

@aussy123 Can you tell me who does that?

Parth (parthkohli):

I'd be happy to know the names of the rule breakers

OpenStudy (anonymous):

lol